OpenStudy (anonymous):

The length of a rectangle is 5 inches longer than its width. The perimeter is 38. Find the dimensions of the rectangle.

Parth (parthkohli):

If the width is \(x\), then the length is \(x + 5\). Use the perimeter formula - \(perimeter = 2(l + w)\)

Parth (parthkohli):

So,\[2(x + x + 5) = 38\]Solve this equation.

OpenStudy (anonymous):

wouldnt it become 2x+2x+10=38?

OpenStudy (anonymous):

and, length=x+5 and breadth =x

OpenStudy (anonymous):

ya

Parth (parthkohli):

No need to make it like that. Divide both sides by 2.

Parth (parthkohli):

\[x + x + 5 = {38 \over 2} \]Etcetera.

OpenStudy (anonymous):

It can be done like that also @herbdog

Parth (parthkohli):

But there's no need to add an extra step :P

OpenStudy (anonymous):

2x+2x+10=38 now subtract 10 from both sides

OpenStudy (anonymous):

becomes 2x+2x=28

OpenStudy (anonymous):

7 by 12 xD

OpenStudy (anonymous):

now 2x+2x =28 4x =28 got it?

OpenStudy (anonymous):

@herbdog

OpenStudy (anonymous):

yes, but dont i divide 4 and 28

OpenStudy (anonymous):

yes now divide both sides by 4 and get the value of x

OpenStudy (anonymous):

you will divide both sides by 4 based on the property of equality. Therefore having x = 7. But the question says about its sides' length so the width is now 7. Add 5 to the width to get the length. 7+5= 12 so length is equal to 12 xD

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

WELCOME